Travel 100 miles in 30 min. What is MPH? With solution please!
Found 2 solutions by jim_thompson5910, unlockmath:
Answer by jim_thompson5910(35256)

30 min = 30/60 hr = 0.5 hr

d = rt

100 = r*0.5

100/0.5 = r

200 = r

r = 200 mph.

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------

Answer by unlockmath(1688)

Hello,
We know that Rate = distance/time
(30 minutes is .5 hour)
Rate=100 mile/.5 hour
Divide and we get:
Rate=200 MPH
(Hint: MPH is Miles per Hour---Per means "for each")
Make sense?
RJ
www.math-unlock.com